What is the length of a second `s pendulum, at a place where acceleration due to gravity is g

Grade:9

1 Answers

Arun
6 years ago
As T = 2π√(l/g) as T = 2 sec so l = g/π² = 9.8/π² = 0.994 m.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer